# VocabTailor: Dynamic Vocabulary Selection for Downstream Tasks in Small Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023a, b; Achiam et al. 2023; Team et al. 2024; Anthropic 2024, 2025; Bai et al. 2023; Guo et al. 2025) have rapidly become foundational to modern AI applications. Recently, increasing attention has turned towards small language models (SLMs), which are better suited for deployment on edge devices and in resource-constrained environments. Despite their compact size, memory still remains a bottleneck, particularly for edge devices with limited GPU memory. A key driver of this bottleneck is the model’s vocabulary size, which directly impacts the memory footprint of both the embedding layer and the language modeling (LM) head. For example, in the Llama 3.2 1B model with a 128K-token vocabulary, the embedding and LM head account for over 20% of the total memory usage. As SLMs are scaled down and deployed under tighter memory constraints, vocabulary-related memory inefficiencies become increasingly unsustainable, posing a fundamental barrier to efficient SLM deployment.
To address this, prior work has explored static vocabulary pruning strategies (Ushio, Zhou, and Camacho-Collados 2023; Yang, Cui, and Chen 2022), which reduce vocabulary size by eliminating rare or irrelevant tokens based on curated corpora. While these approaches are well-motivated, they suffer from key limitations due to their static and coupled design, which assumes a single, globally pruned vocabulary is applied on all vocabulary-related components (i.e.,  tokenizer, embedding, and LM head). This design introduces two major issues:
@@
  • Premature Information Loss: Pruning the tokenizer, embedding, and LM head altogether alters the input representation passed to the transformer. These modified inputs may differ significantly from those seen during pretraining, causing distributional shifts and information loss from the prefill stage. Notably, each pruning step in the pipeline introduces information loss that accumulates, leading to cumulative performance degradation during inference.
  • Lack of flexibility and Adaptability: The static strategy limits the model’s adaptability across diverse tasks. Supporting different task configurations typically requires duplicating multiple copies of vocabulary-related components, resulting in substantial storage overhead and increased complexity in deployment.
Based on empirical observations and theoretical analysis, we derive two key principles for efficient design:
  • Lexical Locality: In common downstream tasks, generation relies on a highly localized vocabulary—each output depends on a small subset of input tokens and a limited set of task-specific tokens.
@@
  • Computation Asymmetry: There is computation asymmetry between embedding and LM head. The embedding layer mainly leverages the lookup operation, which is computationally cheap but limited by memory bandwidth. In contrast, the LM head is compute intensive with massive matrix multiplication requiring immense floating point power, making it an ideal task for a GPU. Existing pruning methods ignore such asymmetry, missing opportunities for system-level optimization.
Guided by these principles, we propose VocabTailor, a flexible and efficient framework for dynamic vocabulary selection. VocabTailor is based on two main pillars:
@@
  • Decoupled Design: We adopt a decoupled design for vocabulary-related components. We retain the full tokenizer and offload the embedding layer to CPU memory. Since the embedding lookup is a memory-intensive operation with $\mathcal{O(1)}$ computational complexity, we can strategically offload it to free up valuable GPU memory, with minimal overhead to the overall system performance.
  • Hybrid Static-Dynamic Vocabulary Selection: At runtime, we dynamically select and load input-relevant tokens while maintaining a small, static set of task-specific tokens to ensure stable and efficient computation in the LM head.
@@
This design enables substantial memory savings without compromising input fidelity and model generality. Compared to static pruning, which retains the union of all input and output tokens $\left(\bigcup\mathcal{I}_{i}\right)\bigcup\left(\bigcup\mathcal{O}_{i}\right)$, VocabTailor only needs $\mathcal{I}_{i}\bigcup\mathcal{T}$ at inference time, where $\mathcal{I}_{i}$, $\mathcal{O}_{i}$ are input and output tokens for example $i$, and $\mathcal{T}$ is a small, fixed task-specific token set. Since $\bigcup\mathcal{I}_{i}\gg\mathcal{I}_{i}$, this leads to substantial memory savings and improved task adaptability.
In summary, this paper makes the following contributions:
@@
    Across five representative downstream tasks—machine translation, summarization, code completion, information extraction, and math problem solving—VocabTailor reduces memory usage of vocabulary-related components by up to 99%, significantly lowers memory usage with minimal or no performance degradation.
@@ 2.1 Small Language Model
Small language models (SLMs) are compact alternatives to large language models (LLMs), which are designed for efficiency, lower computational costs, and deployment on resource-constrained devices. While LLMs like GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023a, b), Claude (Anthropic 2024, 2025), Gemini (Team et al. 2024), Qwen (Bai et al. 2023), and DeepSeek (Guo et al. 2025) have achieved widespread success across real-life applications, SLMs are gaining attention due to their suitability in GPU memory-constrained environments, personal devices, and task-specific scenarios. They offer a scalable, efficient, and sustainable solution tailored for real-time and on-device applications (Lamaakal et al. 2025). With careful selection, small open models can rival and even outperform LLMs while offering improved speed and memory efficiency (Sinha, Jain, and Chadha 2024).
Despite advances in architectures, training techniques, and model compression techniques, SLMs still face challenges, including trade-offs between model size and accuracy, generalization limitations, and concerns over bias and privacy (Van Nguyen et al. 2024; Lamaakal et al. 2025). A common approach to building SLMs is distilling them from LLMs while retaining the same tokenizer and vocabulary. This typically causes vocabulary-related components (i.e.,  embedding and LM Head) to account for a large proportion of the model’s total parameters. This makes vocabulary pruning an effective optimization strategy for SLMs.
@@ 2.2 Tokenization
Tokenization is a fundamental preprocessing step in Natural Language Processing (NLP) that splits text into smaller units called tokens (e.g., words or characters), which form the input to downstream tasks. Over the years, various tokenization techniques have been developed (Sennrich, Haddow, and Birch 2016; Kudo and Richardson 2018; Devlin et al. 2019). Among these, Byte-Pair Encoding (BPE) (Sennrich, Haddow, and Birch 2016) has become one of the most widely used. Originally developed for data compression, BPE was adapted to tokenize text by iteratively merging the most frequent adjacent symbol pairs starting from individual characters until a target vocabulary size is reached. This approach enables efficient representation of frequent words with fewer tokens while breaking rare or unseen words into informative subword units. BPE’s widespread adoption across transformer architectures has established it as a core component of LLM infrastructure.
State-of-the-art LLMs such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023a, b), Gemini (Team et al. 2024), Claude (Anthropic 2024, 2025), and DeepSeek (Guo et al. 2025) rely on BPE-based tokenizers to balance vocabulary efficiency and expressiveness. However, the resulting vocabulary sizes are often large, leading to large embedding matrices and LM heads, which increase computational and memory overhead during inference. This scalability bottleneck has motivated research into vocabulary reduction and adaptive tokenization strategies, especially for resource-constrained deployments.
Importantly, pruning a BPE-based tokenizer can interfere with its learned merge operations governing subword segmentation, resulting suboptimal tokenizations. This disruption reduces the granularity of lexical representations, ultimately impairing model performance on downstream tasks.
@@ 2.3 Vocabulary Pruning
Vocabulary-trimming (VT) (Ushio, Zhou, and Camacho-Collados 2023) and TextPruner (Yang, Cui, and Chen 2022) were proposed to improve efficiency and reduce model size by removing tokens irrelevant to the target language or rarely seen in downstream tasks. Both methods follow a static pruning strategy: they identify language-specific or task-relevant tokens from a curated corpus and prune the vocabulary accordingly. While this process simultaneously reduces the size of the tokenizer, embedding, and LM head, it introduces premature information loss, leading to cumulative performance degradation, and limited flexibility and adaptability, resulting in substantial memory overhead and increased deployment complexity. These two major limitations underscore the need for more flexible and dynamic vocabulary pruning strategies that are task-specific, adaptable, and generalizable across deployment settings. In response, we propose our framework in Section 3.
@@ 3.1 VocabTailor Framework
As illustrated in Fig. 1, VocabTailor decouples vocabulary management across vocabulary-related components. Unlike static vocabulary pruning that uniformly reduces the tokenizer, embedding layer, and LM head, VocabTailor treats each component based on its unique computational and storage properties.
@@
For the embedding layer and LM head, despite their similar roles(tokenize and detokenize), a fundamental computation asymmetry exists between them.
For the embedding layer, which relies on simple lookup operations, is naturally CPU-friendly and incurs negligible runtime overhead when offloaded to CPU memory from high-cost accelerator memory (Yu et al. 2025). Thus, we retain it fully and offload it to CPU.
In contrast, for the LM head, which performs compute-intensive matrix multiplications, must remain on GPU for efficient inference. To optimize its memory footprint, we introduce a hybrid static-dynamic vocabulary selection strategy. This approach leverages the lexical locality of tokens and significantly reduces memory usage while preserving downstream task performance and enabling flexible, task-specific adaptation without model duplication.
@@
Figure 2: Left: Input-output lexical overlap ratio. Right: Example of input-output lexical overlap in summarization task.
@@ 3.2 Motivation: Analysis of Lexical Locality
To investigate the essence of efficient vocabulary selection, we analyze input-output pairs across diverse datasets aligned with various downstream tasks. Our empirical analysis reveals two fundamental properties of lexical locality that motivate VocabTailor’s dynamic vocabulary selection strategy.
@@
In various downstream tasks, the model output reuses tokens from the input, as shown in Fig. 2. This phenomenon is particularly pronounced in text extraction tasks (e.g., span-based QA or named entity recognition), where output tokens are typically a subset of the input tokens. In code completion, generated code often replicates variable names, function names, and other identifiers from the input context. Similarly, in text summarization, the generated summary contains large spans from the source document. For instance, in summarization, on average 61.9% of tokens in generated summaries are copied from the input document. From an information-theoretic perspective, the input context dramatically reduces the entropy of the output token distribution, constraining generation to a much smaller effective vocabulary. Thus, preserving input vocabulary is critical for maintaining performance during vocabulary reduction.
@@
Due to input diversity across datasets, the union of all input tokens is significantly larger than the tokens required for any single generation instance. Each input introduces unique tokens, making the aggregate input vocabulary much larger than individual requirements. Crucially, the remaining output tokens—those not found in the corresponding input—form a relatively small, task-specific set $\mathcal{T}$ that captures the essential generation patterns for the task.
Formally, let $\mathcal{I}$ denote the set of all input tokens, $\mathcal{I}_{i}$ the set of input tokens for instance $i$ in a dataset, $\mathcal{O}$ the set of all output tokens, and $\mathcal{T}\subset\mathcal{O}$ the set of all task-specific tokens in outputs. We observe that:
$\mathcal{I}=\bigcup_{i}\mathcal{I}_{i}\gg\mathcal{I}_{i}\;\text{and }\mathcal{O}>\mathcal{T}$
These observations expose the fundamental inefficiency of static vocabulary pruning that operates on the union $\mathcal{I}\cup\mathcal{O}$ for every example. VocabTailor exploits this lexical locality by dynamically constructing active per-example vocabularies using only per-example input tokens $\mathcal{I}_{i}$ and a compact task-specific set $\mathcal{T}$:
$\mathcal{I}\bigcup\mathcal{O}\gg\mathcal{I}_{i}\bigcup\mathcal{T}$
@@ 3.3 The Hybrid Vocabulary Selection Strategy
Dynamic Selection (Runtime Behavior)
At the start of each inference instance, VocabTailor identifies the unique input token set $\mathcal{I}_{i}$ from the input text and selectively loads the corresponding LM head weight vectors from CPU main memory to GPU. This selective loading exploits the input-driven locality principle, ensuring that only input-relevant vocabulary components are active during generation. Crucially, we overlap this memory transfer with the model’s forward propagation, leveraging GPU-CPU parallelism to mitigate runtime overhead.
Static Selection (Offline Construction)

### v2

@@ Abstract
VocabTailor: Dynamic Vocabulary Selection for Downstream Tasks in Small Language Models
Hanling Zhang1,2thanks: Equal contribution., Yayu Zhou611footnotemark: 1, Tongcheng Fang 3, Zhihang Yuan3thanks: Co-corresponding author., Guohao Dai 1,5, Wanli Ouyang 2,4, Yu Wang 322footnotemark: 2 1Infinigence AI, 2The Chinese University of Hong Kong, 3Tsinghua University, 4Shanghai Artificial Intelligence Laboratory, 5Shanghai Jiao Tong University, 6Independent Researcher
@@ 1 Introduction
Large Language Models (LLMs) (brown2020language; touvron2023llama; touvron2023llama-1; achiam2023gpt; team2024gemini; anthropic_claude_2024; anthropic_introducing_2025; bai2023qwen; guo2025deepseek) have rapidly become foundational to modern AI applications. Recently, increasing attention has turned towards small language models (SLMs), which are better suited for deployment on edge devices and in resource-constrained environments. Despite their compact size, memory still remains a bottleneck, particularly for edge devices with limited GPU memory. A key driver of this bottleneck is the model’s vocabulary size, which directly impacts the memory footprint of both the embedding layer and the language modeling (LM) head. For example, in the Llama 3.2 1B model with a 128K-token vocabulary, the embedding and LM head account for over 20% of the total memory usage. As SLMs are scaled down and deployed under tighter memory constraints, vocabulary-related memory inefficiencies become increasingly unsustainable, posing a fundamental barrier to efficient SLM deployment.
To address this, prior work has explored static vocabulary pruning strategies (ushio_efficient_2023; yang_textpruner_2022), which reduce vocabulary size by eliminating rare or irrelevant tokens based on curated corpora. While these approaches are well-motivated, they suffer from key limitations due to their static and coupled design, which assumes a single, globally pruned vocabulary is applied on all vocabulary-related components (i.e.,  tokenizer, embedding, and LM head). This design introduces two major issues. First, premature information loss arises because pruning the tokenizer, embedding, and LM head altogether alters the input representation passed to the transformer. These modified inputs may differ significantly from those seen during pretraining, causing distributional shifts and information loss from the prefill stage. Notably, each pruning step in the pipeline introduces information loss that accumulates, leading to cumulative performance degradation during inference. Second, lack of flexibility and adaptability as the static strategy limits the model’s adaptability across diverse tasks. Supporting different task configurations typically requires duplicating multiple copies of vocabulary-related components, resulting in substantial storage overhead and increased complexity in deployment.
@@
Based on empirical observations and theoretical analysis, we derive two key principles for efficient design. Lexical locality captures the empirical observations that, in common downstream tasks, generation relies on a highly localized vocabulary, where each output depends on a small subset of input tokens and a limited set of task-specific tokens. Computation asymmetry reflects the distinct computational characteristics of the embedding and LM head: the embedding layer mainly leverages lookup operations, which are computationally cheap but memory-bandwidth bound, whereas the LM head is compute-intensive with massive matrix multiplication requiring immense floating-point power and thus better suited for GPUs. Existing pruning methods ignore such asymmetry, missing opportunities for system-level optimization.
@@
Guided by these principles, we propose VocabTailor, a flexible and efficient framework for dynamic vocabulary selection. VocabTailor is based on two main pillars. First, we adopt a decoupled design for vocabulary-related components. We retain the full tokenizer and offload the embedding layer to CPU memory. Since the embedding lookup is a memory-intensive operation with $\mathcal{O}(1)$ computational complexity, we can strategically offload it to free up valuable GPU memory, with minimal overhead to the overall system performance. Second, we propose hybrid static-dynamic vocabulary selection: at runtime, we dynamically select and load input-relevant tokens while maintaining a small, static set of task-specific tokens to ensure stable and efficient computation in the LM head.
@@
This design enables substantial memory savings without compromising input fidelity and model generality. Compared to static pruning, which retains the union of all input and output tokens $\left(\bigcup\mathcal{I}_{i}\right)\bigcup\left(\bigcup\mathcal{O}_{i}\right)$, VocabTailor only needs $\mathcal{I}_{i}\bigcup\mathcal{T}$ at inference time, where $\mathcal{I}_{i}$, $\mathcal{O}_{i}$ are input and output tokens for example $i$, and $\mathcal{T}$ is a small, fixed task-specific token set. Since $|\bigcup\mathcal{I}_{i}|\gg|\mathcal{I}_{i}|$, this leads to substantial memory savings and improved task adaptability.
In summary, our contributions are:
@@
    Across five representative downstream tasks—machine translation, summarization, code completion, information extraction, and math problem solving—VocabTailor reduces memory usage of vocabulary-related components by up to 99%, with minimal or no performance degradation.
@@ 2.1 Small Language Model
Small language models (SLMs) are compact alternatives to large language models (LLMs), which are designed for efficiency, lower computational costs, and deployment on resource-constrained devices. While LLMs like GPT-4 (achiam2023gpt), LLaMA (touvron2023llama; touvron2023llama-1), Claude (anthropic_claude_2024; anthropic_introducing_2025), Gemini (team2024gemini), Qwen (bai2023qwen), and DeepSeek (guo2025deepseek) have achieved widespread success across real-life applications, SLMs are gaining attention due to their suitability in GPU memory-constrained environments, personal devices, and task-specific scenarios. They offer a scalable, efficient, and sustainable solution tailored for real-time and on-device applications (lamaakal2025tiny). With careful selection, small open models can rival and even outperform LLMs while offering improved speed and memory efficiency (sinha2024small).
Despite advances in architectures, training techniques, and model compression techniques, SLMs still face challenges, including trade-offs between model size and accuracy, generalization limitations, and concerns over bias and privacy (van2024survey; lamaakal2025tiny). A common approach to building SLMs is distilling them from LLMs while retaining the same tokenizer and vocabulary. This typically causes vocabulary-related components (i.e.,  embedding and LM Head) to account for a large proportion of the model’s total parameters. This makes vocabulary pruning an effective optimization strategy for SLMs.
@@ 2.2 Tokenization
Tokenization is a fundamental preprocessing step in Natural Language Processing (NLP) that splits text into smaller units called tokens (e.g., words or characters), which form the input to downstream tasks. Over the years, various tokenization techniques have been developed (sennrich_neural_2016; kudo2018sentencepiece; devlin_bert_2019). Among these, Byte-Pair Encoding (BPE) (sennrich_neural_2016) has become one of the most widely used. Originally developed for data compression, BPE was adapted to tokenize text by iteratively merging the most frequent adjacent symbol pairs starting from individual characters until a target vocabulary size is reached. This approach enables efficient representation of frequent words with fewer tokens while breaking rare or unseen words into informative subword units. BPE’s widespread adoption across transformer architectures has established it as a core component of LLM infrastructure.
State-of-the-art LLMs such as GPT-4 (achiam2023gpt), LLaMA (touvron2023llama; touvron2023llama-1), Gemini (team2024gemini), Claude (anthropic_claude_2024; anthropic_introducing_2025), and DeepSeek (guo2025deepseek) rely on BPE-based tokenizers to balance vocabulary efficiency and expressiveness. However, the resulting vocabulary sizes are often large, leading to large embedding matrices and LM heads, which increase computational and memory overhead during inference. This scalability bottleneck has motivated research into vocabulary reduction and adaptive tokenization strategies, especially for resource-constrained deployments.
@@ 2.3 Vocabulary Pruning
Vocabulary-trimming (ushio_efficient_2023) and TextPruner (yang_textpruner_2022) were proposed to improve efficiency and reduce model size by removing tokens irrelevant to the target language or rarely seen in downstream tasks. Both methods follow a static pruning strategy: they identify language-specific or task-relevant tokens from a curated corpus and prune the vocabulary accordingly. While this process simultaneously reduces the size of the tokenizer, embedding, and LM head, it introduces premature information loss, leading to cumulative performance degradation, and limited flexibility and adaptability, resulting in substantial memory overhead and increased deployment complexity. These two major limitations underscore the need for more flexible and dynamic vocabulary pruning strategies that are task-specific, adaptable, and generalizable across deployment settings. In response, we propose our framework in Section 3.
@@ 3.1 VocabTailor Framework
VocabTailor (shown in Figure 1) decouples vocabulary management across vocabulary-related components. Unlike static vocabulary pruning that uniformly reduces the tokenizer, embedding layer, and LM head, VocabTailor treats each component based on its unique computational and storage properties.
@@
For the embedding layer and LM head, despite their similar roles(tokenize and detokenize), a fundamental computation asymmetry exists between them. The embedding layer, which relies on simple lookup operations, is naturally CPU-friendly and incurs negligible runtime overhead when offloaded to CPU memory from high-cost accelerator memory yu_scaling_2025. Thus, we retain it fully and offload it to CPU. In resource-constrained deployment settings such as edge devices or platforms with unified CPU–GPU memory, even CPU-resident embeddings may impose non-trivial memory overhead. To address this, VocabTailor supports disk-backed embedding offloading using the Lightning Memory-Mapped Database (LMDB). Embedding vectors are stored as key–value entries and retrieved on demand during inference, preserving full-vocabulary access while further reducing memory footprint. Implementation details and analysis are in Appendix B.
The LM head, which performs compute-intensive matrix multiplications, must remain on GPU for efficient inference. To optimize its memory footprint, we introduce a hybrid static-dynamic vocabulary selection strategy. This approach leverages the lexical locality and significantly reduces memory usage while preserving downstream task performance and enabling flexible, task-specific adaptation without model duplication.
@@ 3.2 Motivation: Analysis of Lexical Locality
To investigate the essence of efficient vocabulary selection, we analyze input-output pairs across diverse datasets aligned with downstream tasks. Our empirical analysis reveals two fundamental properties of lexical locality that motivate VocabTailor’s dynamic vocabulary selection strategy.
@@
Figure 2: Left: Input-output lexical overlap ratio. Right: Example of lexical overlap in a summarization task.
In various downstream tasks, the model output reuses tokens from the input (shown in Figure 2). This phenomenon is particularly pronounced in text extraction tasks (e.g., span-based QA or named entity recognition), where output tokens are typically a subset of the input tokens. In code completion, generated code often replicates variable names, function names, and other identifiers from the input context. Similarly, in text summarization, the generated summary contains large spans from the source document. For instance, in summarization, on average 61.9% of tokens in generated summaries are copied from the input document. From an information-theoretic perspective, the input context dramatically reduces the entropy of the output token distribution, constraining generation to a much smaller effective vocabulary. Thus, preserving input vocabulary is critical for maintaining performance during vocabulary reduction.
@@
Due to input diversity across datasets, the union of all input tokens is significantly larger than the tokens required for any single generation instance. Each input introduces unique tokens, making the aggregate input vocabulary much larger than individual requirements. The remaining output tokens—those not found in the corresponding input—form a relatively small, task-specific set $\mathcal{T}$ that captures the essential generation patterns for the task.
Formally, let $\mathcal{I}$ denote the set of all input tokens, $\mathcal{I}_{i}$ the set of input tokens for instance $i$ in a dataset, $\mathcal{O}$ the set of all output tokens, and $\mathcal{T}\subset\mathcal{O}$ the set of all task-specific tokens in outputs. We observe:
$|\mathcal{I}|=\left|\bigcup_{i}\mathcal{I}_{i}\right|\gg|\mathcal{I}_{i}|\;\text{and }|\mathcal{O}|>|\mathcal{T}|$
These observations expose the fundamental inefficiency of static vocabulary pruning hat operates on the union $\mathcal{I}\cup\mathcal{O}$ for every example. VocabTailor exploits this lexical locality by dynamically constructing active per-example vocabularies using only per-example input tokens $\mathcal{I}_{i}$ and a compact task-specific set $\mathcal{T}$:
$\left|\mathcal{I}\bigcup\mathcal{O}\right|\gg\left|\mathcal{I}_{i}\bigcup\mathcal{T}\right|$
@@ 3.3 The Hybrid Vocabulary Selection Strategy
3.3.1 Dynamic Selection (Runtime Behavior)
At the start of each inference instance, VocabTailor identifies the unique input token set $\mathcal{I}_{i}$ from the input text and selectively loads the corresponding LM head weight vectors from CPU to GPU. This selective loading exploits the input-driven locality principle, ensuring that only input-relevant vocabulary components are active during generation.
Beyond selecting input-related tokens at runtime, VocabTailor must efficiently extend the LM head on the GPU. A naive implementation that repeatedly concatenates static and dynamic weights and instantiates new linear layers incurs unnecessary CPU–GPU transfers and module creation overhead. To mitigate this, VocabTailor utilizes the characteristics of input-related tokens and implements a pre-allocated buffer to amortize LM head construction across inputs. Implementations and performance evaluation are detailed in Appendix A.
3.3.2 Static Selection (Offline Construction)